OpenStudy (anonymous):

In five years, Josh will be 21 years old. How old is he now? Let J = Josh's current age. Which formula would calculate Josh's current age?

OpenStudy (anonymous):

J + 5 = 21 j = 21-5 J = 16

OpenStudy (anonymous):

it was the first one j+5=21

OpenStudy (anonymous):

The formula is the first one the second one is how to continue on that formula to find his age

OpenStudy (anonymous):

k